# Rates of Convergence of Generalised Variational Inference Posteriors under Prior Misspecification

### v1

@@ Abstract
We prove rates of convergence and robustness to prior misspecification within a Generalised Variational Inference (GVI) framework with bounded divergences. This addresses a significant open challenge for GVI and Federated GVI that employ a different divergence to the Kullback–Leibler under prior misspecification, operate within a subset of possible probability measures, and result in intractable posteriors. Our theoretical contributions cover severe prior misspecification while relying on our ability to restrict the space of possible GVI posterior measures, and infer properties based on this space. In particular, we are able to establish sufficient conditions for existence and uniqueness of GVI posteriors on arbitrary Polish spaces, prove that the GVI posterior measure concentrates on a neighbourhood of loss minimisers, and extend this to rates of convergence regardless of the prior measure.
Rates of Convergence,
keywords:
[class=MSC]
keywords:
@@
, , and
1 Introduction
The Bayesian, by nature of being Bayesian, can elicit arbitrary priors and may hence encounter prior misspecification; she must either suppose the prior encodes the best available knowledge of the underlying process—this however is scarcely practicable and processes exist for which we cannot intuit priors that lead to consistent posteriors—or accept the arbitrariness [14, 22]. In fact, for any finite number of observations the prior alone may induce any measure a posteriori. Moreover, it is scarcely the case that the model is well specified, that is the model family $\Theta$ will not contain the data generating process [7, 31, 34]. Generalised Variational Inference (GVI) [24] recasts Bayesian posterior updating as an optimisation problem in the calculus of variation over a simpler space of measures, allowing decision makers to mitigate misspecification of both the model and the prior.
The present paper is concerned with characterising the dynamics of updating a posterior measure in the presence of prior and model misspecification through Generalised Variational Inference, as introduced by Knoblauch, Jewson and Damoulas [24]. In particular, we aim to answer the following open problems: under what conditions is a Generalised Variational Inference posterior measure robust to prior misspecification? Do GVI posteriors exist when observations come from an infinite dimensional stochastic process and are these asymptotically consistent? At what rate does this convergence take place?
Asymptotic consistency of posterior measures justifies the updating scheme over its frequentist counterpart by illustrating that we may come arbitrarily close to the optimal given sufficient observations. In Bayesian inference under misspecification asymptotic consistency is often credited when disregarding model and prior misspecification [22, 31, 21]. In fact, Bayesian posteriors concentrate around hypothesis measures in the support of the prior that minimise the Kullback–Leibler divergence rate to the data generating process [30, 7, 5, 15, 32, 33, 22, 16, 31, 34]. This, however, embeds several assumptions which may lead to undesirable posteriors in the misspecified setting: (a) the Kullback–Leibler rate as a choice of loss function, (b) the prior is well specified in that it does not discard hypotheses desired by the decision maker, and leads to consistent posteriors (see e.g. [13, 14]), and moreover, (c) that Bayes is computationally tractable, and hence of use to decision makers.
We already established that (b) may not hold as priors that do not support neighbourhoods around hypotheses with minimal KL divergence rate to the data generating process will lead to concentration around less favourable hypotheses, especially in well specified models, not to mention natural choices of priors that lead to inconsistent posteriors as observed in [13]. Furthermore, since the Kullback–Leibler divergence is known to not be robust (cf. [18]), (a) may lead to suboptimal performance. And as Bayesian inference is rarely computationally tractable, optimising over a simpler space of probability measures—as done in Variational Inference with e.g. the set of Gaussians [19]—allows for computational feasibility. Generalised Variational Inference [24] allows us to mitigate (a), (b), and (c).
[…]
Figure 1: A taxonomy of Generalised Variational Inference. We highlight some special cases.
Suppose $X_{1},X_{2},$... is a sequence of random variables distributed according to some unknown, infinite dimensional data generating process $P_{0}$, and which are modelled through hypothesis measures in the hypothesis space $\Theta$. We begin with some prior measure $\Pi$ and after receiving observations we may update this to $Q_{1},Q_{2},$... accordingly. GVI assumes an optimisation centric viewpoint where we may mitigate (a) by replacing the average negative log–likelihood by an arbitrary loss—such as those based on robust divergences between the data generating process and the hypothesis measure—(b) by replacing the Kullback–Leibler divergence to the prior by an arbitrary divergence, and (c) by optimising over a ‘simpler’ space of measures; see Figure 1 for details. We therefore aim to find a measure within a subset of possible Borel probability measures, $\mathcal{Q}\subset\mathcal{P}(\Theta)$, that achieves
$\inf_{Q\in\mathcal{Q}}\left\{\mathbb{E}_{Q}[n\cdot L(X_{1}^{n},\,\theta)]+\frac{1}{\beta}D(Q:\Pi)\right\}\,.$ (1)
As we are concerned with minimising this functional over the set of possible probability measures, we may write this as
$T_{n}(Q):=n\cdot J_{L_{n}}(Q)+\frac{1}{\beta}D(Q:\Pi)$ (2)
where $J_{L_{n}}(\cdot)$ is the functional given by integration of the loss with $n$ data points against some probability measure. Our results do not require there to be any data, and we may simply drop this dependence whenever convenient. The GVI posterior measure, $Q_{n}$ then minimises this
$Q_{n}\in\operatorname*{arg\,min}_{Q\in\mathcal{Q}}\,T_{n}(Q)$ (3)
whenever it exists. This framework allows decision makers to target a broader set of possible posterior measures, depending on the choices of loss function and divergence, rather than the standard Bayesian posterior; in the case where no misspecification occurs—the prior and likelihood are well specified—the Bayesian posterior is indeed optimal [23, 24, 26, 27]. See Figure 1 for a taxonomy of the GVI framework relating (a), (b), and (c) with the corresponding solutions.
Related Theoretical Work
Rates of convergence of Bayesian posteriors under model misspecification, mild prior misspecification, and highly correlated data are well understood [7, 22, 16, 31]. For generalised Bayesian inference, asymptotic consistency, but not rates of convergence, has been established by Miller [28] in fairly general settings. These concentration results do not apply under severe prior misspecification, requiring the prior to have positive measure on every Kullback–Leibler neighbourhood of the underlying model. Knoblauch [23] established early results on asymptotic consistency of GVI in the parametric setting under fairly strong assumptions, most of which we are able to significantly relax while achieving stronger results.
Main Contributions
@@ 1 Introduction
This paper develops theory concerning GVI posterior measures over infinite dimensional hypothesis spaces $\Theta$. The main contribution of this paper is Theorem 7 where we establish rates of convergence of GVI posterior measures just slower that $n^{-1}$ when using bounded divergences. In doing so we first present sufficient conditions for existence and uniqueness of GVI posteriors, which we then use as a basis to characterise asymptotic concentration to sets containing the minimisers of the loss. Significantly, our theory allows us to deal with instances in which the prior is severely misspecified, such as in the case where the prior has full measure on a subset of $\Theta$ where the Kullback–Leibler divergence rate is infinite and where the prior and GVI posterior are mutually singular. This is highlighted in Figure 1.
Organisation
@@
In Section 2 we provide the necessary background for Generalised Variational Inference, prior misspecification, and the assumptions used throughout the paper. Section 3 first states and discusses the main results of this paper and we provide the complete proofs of these in Section 5. To highlight our results, we present illustrative examples in Section 4 comparing GVI with Bayesian Inference under prior misspecification in Gaussian location models. In Section 6 we provide some concluding remarks on our results, their relevance to extensions of GVI in Federated Learning [27], and relevant open problems.
2 Preliminaries on Generalised Variational Inference and Prior Misspecification
@@
Let $X_{1},X_{2},...,X_{n}$ be a sequence of observable random variables, short $X_{1}^{n}$, generated by the abstract probability space $(\Omega,\mathcal{F},P_{0})$ and taking values in the measurable space $(\Xi,\mathcal{X})$ with $P_{0}$ distribution. We consider $n\in\mathbb{N}$ and let this tend to infinity for the concentration results. Denote by $\sigma(X_{1}^{n})$ the natural filtration of this sequence. We do not place any restrictions on the data, as long as the loss associated with it satisfies the respective assumptions, especially Assumption 7.
@@
Throughout, we will study probability measures in $\mathcal{P}(\Theta),$ the space of all Borel probability measures on $(\Theta,\mathcal{T})$ [20, 11]. We begin with some prior $\Pi$ and update this to the GVI posterior $Q_{n}$.
@@
2.1 Generalised Variational Inference and Related Methods
The Bayesian posterior—consisting of the likelihood function of the data, $p_{\theta}(X_{1}^{n})$, and a prior measure, $\Pi$—is
$\Pi_{n}A:=\cfrac{\int_{A}p_{\theta}(X_{1}^{n})\,\Pi(d\theta)}{\int_{\Theta}p_{\theta}(X_{1}^{n})\,\Pi(d\theta)}$
as usual. If we replace the likelihood by the exponential of a negative loss acting on the data we get Generalised Bayesian Inference (GBI) [9, 28], or a Gibbs posterior [2]. If we discard the prior, we may instead get an energy–based, frequentist model. As shown in [36] Bayes may be equivalently formulated as an optimisation problem where $\Pi_{n}$ is the posterior minimising the ‘Evidence Lower Bound’—the expectation of the negative log–likelihood, and the Kullback–Leibler divergence to the prior—over all probability measures acting on the hypothesis space. GVI instead replaces the negative log–likelihood by any loss function, similar to GBI, as well as the KL divergence by any divergence, generalising these previous approaches. Tangentially, minimising a divergence other than the KL divergence to the Bayesian posterior, not just the prior, may result in nicer optimisation problems in practice, e.g. in Rényi variational inference [25]. This however does not nicely decompose into an expectation acting on the loss and a divergence measure acting on the prior, conflating model and prior misspecification; see also section 2.3.2 in [24]. Typical choices of loss functions are those that either offer attractive computational benefits, and those that incorporate the hypothesis measure robustly, including those based on robust divergences [27].
2.2 Prior Misspecification
As identified in [34, 31] the elements of $\Theta$ that the Bayesian posterior targets are those minimising the Kullback–Leibler divergence between the data generating process and the hypothesis measure. Denoting this as a function on $\Theta$, we write:
$h(\theta)=\lim_{n}\frac{1}{n}\mathbb{E}_{P_{0}}\left[\log\frac{p_{0}(x_{1}^{n})}{p_{\theta}(x_{1}^{n})}\right]$
which may possibly be infinite; note that the expectation is with respect to $P_{0}$ and hence over $\Omega$. This is the loss function targeted in classical Bayesian inference, in particular the elements of $\Theta$ that minimise the above minimise the Kullback–Leibler divergence between the data generating process and the hypothesis measure. If the prior has full measure on the set of elements where this rate is infinite, denote this as $I:=\{\theta\in\Theta:h(\theta)=\infty\}$, then all previous results on consistency, let alone rates of convergence, of (Generalised) Bayesian inference fail [31]. In this case we say that the prior is severely misspecified. In this we also include priors that lead to inconsistent posteriors such as those observed in [13].
To make this rigorous, we place the following definition of robustness to prior misspecification.
Definition 1.
Classical results on Bayesian sensitivity analysis deal with small perturbations of some reference prior by some other measure through additive or multiplicative misspecification [12, 17]. That is some reference prior $\Pi$ is perturbed by another probability measure $G\in\mathcal{P}(\Theta)$ through a constant $\varepsilon\in[0,1]$, e.g. $\Pi_{\varepsilon}=(1-\varepsilon)\Pi+\varepsilon G$. Our results are different to this line of work, and we will not use such explicit formulations for $\mathcal{G}$. We discuss this distinction in Appendix B.
2.3 Illustrative Example: Bayesian Inference under Prior Misspecification
To illustrate the non–robustness of Bayesian inference under prior misspecification, we give a simple, parametric example with $\Theta=\mathbb{R}$. Consider the classical Bayesian posterior, and assume that $p_{0}$ is Gaussian with unknown mean and known variance, $X_{i}\sim\mathcal{N}(\theta_{0},\sigma_{p}^{2})$. The prior is a contaminated Gaussian, $\pi\sim\mathcal{N}(\mu_{\pi},\sigma_{\pi}^{2})$ where $\mu_{\pi}\in\mathbb{R}$. Then the posterior on the mean, $\theta$, is Gaussian and characterised by the following variance and mean:
$\sigma_{n}^{2}=\left(\sigma_{\pi}^{-2}+n\sigma_{p}^{-2}\right)^{-1},\,$ $\mu_{n}=\sigma_{n}^{2}\left(\frac{\mu_{\pi}}{\sigma_{\pi}^{2}}+\sum_{i=1}^{n}\frac{X_{i}}{\sigma_{p}^{2}}\right)$
For fixed $n$, and given knowledge of the data, we may pick $\mu_{\pi}$ and $\sigma_{\pi}$ to achieve any Gaussian distribution a posteriori, disregarding the data entirely.
Moreover, the Kullback–Leibler divergence to the prior has closed form:
$D_{KL}(q:\pi)=\log\frac{\sigma_{\pi}}{\sigma_{q}}+\frac{(\mu_{q}-\mu_{\pi})^{2}+\sigma_{q}^{2}}{2\sigma_{\pi}^{2}}-\frac{1}{2}$
which is convex in both $\mu_{q}$ and $\sigma_{q}$. Therefore, as $\mu_{\pi}\rightarrow\infty$, this increases, and hence forces $\mu_{q}$ to match the prior mean as the gradient is minimised if and only if their means are equal. See Appendix C for further details.
2.4 Assumptions
For the GVI objective to make sense we require some standard assumptions that will ensure the problem is not pathological.
@@ 2.4 Assumptions
The losses $L(X_{1}^{n},\,\theta)$ are $\sigma(X_{1}^{n})\times\mathcal{T}$–measurable and lower bounded [^2] Loss functions used in ML are typically non–negative, we, in accordance with the literature on GVI and GBI, allow for more general functions through the lower boundedness. for any $n\in\mathbb{N}$.
This standard assumption ensures that the GVI objective is not pathological, making sense as a machine learning objective. Allowing the loss to be $\infty$’ would imply that we could pick a Dirac–delta measure at any such point. However, we do not place any restrictions on the underlying data distribution, and strictly speaking, our results do not require there to be any data at all; we could simply work with a sequence of $\mathcal{T}$–measurable functions as in [28]. Where this distinction is not necessary, to lighten notation, we may write $L_{n}(\theta):=L(X_{1}^{n},\,\theta)$.
Furthermore, we require that the subspace of “simpler” measures we are optimising over is well behaved and in particular that we are able to find an element of this space for which the GVI functional is finite.
@@
Further assumptions on this space $\mathcal{Q}$ such as it being closed in the weak topology or being convex are made explicit in the theorems where required. These are sufficient to ensure existence and uniqueness of GVI posteriors as soon as there exists a single element of $\mathcal{Q}$ for which Equation 2 is finite (see Theorems 2, 4 and 8).
As we desire a framework that is able to deal with severe prior misspecification, we require the following assumption on the divergence.
@@
The assumption here is a stricter requirement and violated by many common divergences, such as the KL divergence, as shown in the previous section. However such bounded–above divergences do exist, a prime example being the Total Variation Distance. This assumption allows us to bound the influence of the prior on the GVI objective, and forces the GVI posterior to not discard the data. We find that this allows for favourable behaviour of GVI in the presence of prior misspecification, ensuring frequentist consistency under very general assumptions, including the case where the GVI posterior and the prior are mutually singular.
@@
It is worth noting that this may be satisfied by both bounded divergences, such as the total variation distance, as well as specific choices of the set of measures we are optimising over, $\mathcal{Q}$, and the set of possible priors, $\mathcal{G}$. In the former case, we may allow for $\mathcal{G}=\mathcal{P}(\Theta)$ and $Q\subset\mathcal{P}(\Theta)$ without restriction. The latter is more restrictive and includes the case where $\mathcal{G}$ is a singleton and $Q$ the sublevel sets of $\mathcal{P}(\Theta)$.
The next assumption is of a technical nature, saying that we may achieve the infimum of the loss functional over the set $\mathcal{Q}$ with some element from that set.
Assumption 4.
$\exists P_{n}^{\star}\in\mathcal{Q}$ such that $\forall Q\in\mathcal{Q}$ it holds that $J_{L_{n}}(P_{n}^{\star})\leq J_{L_{n}}(Q)$.
This is not a particularly restrictive assumption, and will generally be satisfied by typical choices of $\mathcal{Q}$. We use the ‘’ notation throughout to denote minimisers, or a relation to a minimiser.
To prove existence and uniqueness of the GVI posterior, we require additional assumptions on the loss and the divergence.
Assumption 5.
The losses as functions on $\Theta$, $L_{n}:\Theta\rightarrow\mathbb{R}$, are coercive and lower semi–continuous for all $n\in\mathbb{N}$.
This assumption is crucial to establish the existence of minimisers as we require coerciveness of the loss in order to establish tightness of a sequence of probability measures. The lower semi–continuity of the loss is easier to verify and the same then immediately follows for the functional [11, 4], however it is slightly stronger than required as we merely need the lower semi–continuity of the functional $J_{L_{n}}(\cdot)$.
Remark.
This assumption carries several implications. The lower semi–continuity of the loss allows it to achieve its infima. If $\mathcal{Q}$ is weak closed then we may discard Assumption 4 as such a measure will necessarily exist. Note also that the coercivity of the loss may be discarded if $\Theta$ were compact.
Assumption 6.
The divergence $D:\mathcal{P}(\Theta)\times\mathcal{P}(\Theta)\rightarrow\mathbb{R}_{\geq 0}$ is lower semi–continuous in its first argument.
In combination with Assumption 5, this establishes that the entire GVI functional is lower semi–continuous, a key tool in proving the existence of minimisers to Equation 3, and is satisfied by key divergences, such as the KL divergence or the total variation distance.
Finally, in order to reason about the asymptotic properties and rates of convergence of the GVI posteriors we need that:
@@
This assumption is required for the concentration results of Theorems 6 and 7, and is standard in Bayesian asymptotics (see e.g. [31]) where an asymptotic equipartition property, as in the Shannon–MacMillan–Breimann theorem, is required to hold. Unless otherwise stated all convergence takes place in the number of observables $n$.

### v2

@@ Abstract.
We prove rates of convergence and robustness to prior misspecification within a Generalised Variational Inference (GVI) framework with bounded divergences. This addresses a significant open challenge for GVI and Federated GVI that employ a different divergence to the Kullback–Leibler under prior misspecification, operate within a subset of possible probability measures, and result in intractable posteriors. Our theoretical contributions extend to misspecified priors that lead to inconsistent Bayes posteriors. In particular, we are able to establish sufficient conditions for existence and uniqueness of GVI posteriors on arbitrary Polish spaces, prove that the GVI posterior measure concentrates on a neighbourhood of loss minimisers, and extend this to rates of convergence regardless of the prior measure.
Key words and phrases:
Generalised Variational Inference, Prior Misspecification, Model Misspecification, Frequentist Consistency, Rates of Convergence.
2020 Mathematics Subject Classification:
Primary 62G20; secondary 62G35, 62C10, 62F12, 62F15, 62F35.
1. Introduction
Consider the problem of posterior inference after having constructed a prior measure $\Pi$ and a hypothesis space $(\Theta,\mathcal{T})$ with associated hypothesis measures in the model ${\mathcal{P}}=\{P_{\theta}:\theta\in\Theta\}$. We are concerned with situations in which (a) the prior is misspecified, leading to inconsistent Bayes posteriors [diaconis1986a, diaconis1986b], (b) the model is misspecified, not containing the data generating process [kleijn2006], and (c) computational capabilities limit our ability to conduct posterior inference [pinski2015].
To this effect we consider posterior inference through Generalised Variational Inference (GVI) as introduced by Knoblauch, Jewson, and Damoulas [jeremias2022]. The present paper is concerned with characterising the dynamics of updating a posterior measure in the presence of (a), (b), and (c) through GVI and places particular emphasis on prior misspecification. In particular, we aim to answer the following open problems: under what conditions is a Generalised Variational Inference posterior measure robust to prior misspecification? Do GVI posteriors exist when observations come from an infinite dimensional stochastic process and are these asymptotically consistent? At what rate does this convergence take place?
Throughout, we consider $X_{1},X_{2},$..., taking values in $(\Xi,\mathcal{X})$, a sequence of observable random variables with unspecified dependencies generated by some unknown data generating process with $P_{0}$ its infinite dimensional process distribution. GVI recasts Bayesian updating as a problem in the calculus of variation, where we link observables and hypotheses through some loss function $L(X_{1}^{n},\,\theta)$, regularise with respect to the prior through a divergence $D(\cdot:\Pi)$, and optimise over a restricted space of possible posterior measures ${\mathcal{Q}}\subset\mathcal{P}(\Theta)$. The GVI posterior is defined as
(1) $Q_{n}\in\operatorname*{arg\,min}_{\nu\in\mathcal{Q}}\,\left\{T_{n}(\nu):=n\cdot J_{L_{n}}(\nu)+\beta^{-1}D(\nu:\Pi)\right\}$
whenever it exists. Here $J_{L_{n}}(\nu):={\mathbb{E}}_{\nu}[L(X_{1}^{n},\,\theta)]$ is the linear functional given by integration of the loss with $n$ observables against some probability measure $\nu\in{\mathcal{Q}}$ and the constant $\beta>0$ denotes a learning rate parameter. As our results do not require there to be any data, we may simply drop this dependence whenever convenient and denote $L(X_{1}^{n},\theta)=L_{n}(\theta)$.
The study of asymptotic consistency is classical for Bayes as it justifies an updating scheme over its frequentist counterpart by illustrating that we may come arbitrarily close to the optimal given sufficient observations [schwartz1965, berk1966, diaconis1986a, barron1999, ghosal2000, shen2001, walker2004, ghosal2007, shalizi2009, walker2013]. Considering now the question of asymptotic consistency of posterior measures reveals why it is advantageous to study GVI posteriors rather than their (Generalised) Bayesian counterparts under prior misspecification. Consider $(\Pi_{n})_{n}$ the sequence of Bayes posteriors, and $d(\cdot,\cdot)$ a dissimilarity score between hypothesis measures, then the Bayes posteriors are consistent if for all $\varepsilon>0$ it holds with $P_{0}$ probability 1 that
$\Pi_{n}(P_{\theta}\in{\mathcal{P}}\cap\operatorname*{supp}{\Pi}:d(P_{\theta},P_{0})>\varepsilon\,|\,X_{1},X_{2},...,X_{n})\rightarrow 0.$
The choice of dissimilarity measure $d(\cdot,P_{0})$ now illustrates our interest in GVI. A typical such measure is given as $d(P_{\theta}:P_{0})=D_{KL}(P_{\theta}:P_{0})-\operatorname*{ess\,inf}_{Q\in{\mathcal{P}}}^{\Pi}D_{KL}(Q:P_{0})$ where the essential infimum over $\mathcal{P}$} is taken with respect to the prior $\Pi$, see for instance [berk1966, walker2004, shalizi2009]. Further, we may only concentrate on measures in the support of the prior which is immediate in the formulation of the Bayes posterior or its variational counterpart in Equation 1 (see [zellner1988]).
The Bayesian, by nature of being Bayesian, can elicit arbitrary priors and may hence encounter prior misspecification; she must either suppose the prior encodes the best available knowledge of the underlying process—this however is scarcely practicable and processes exist for which we cannot intuit priors that lead to consistent posteriors—or accept the arbitrariness [diaconis1986b, kleijn2006]. In fact, the nonparametric setting intrinsically encumbers the choice of reasonable prior as natural choices may lead to inconsistency [diaconis1986a, diaconis1986b], due to these not satisfying structural assumptions as required for consistency [kleijn2006, shalizi2009]. Hence, the prior may be misspecified.
Moreover, it is scarcely the case that the model is well specified, that is the model family $\Theta$ will not contain the data generating process [berk1966, shalizi2009, walker2013]. Furthermore, since the Kullback–Leibler divergence is known to not be robust (cf. [jewson2018]), (b) may lead to suboptimal performance due to the choice of the negative log likelihood as a loss function. Lastly, as Bayesian inference is rarely computationally tractable, optimising over a simpler space of probability measures—as done in Variational Inference with e.g. the set of Gaussians [katsevich2024]—allows for computational feasibility.
Generalised Variational Inference recasts Bayesian posterior updating as an optimisation problem in the calculus of variation over a simpler space of measures, allowing decision makers to mitigate (a), (b), and (c); see Figure 1 for reference.
@@ 1. Introduction
Figure 1. A taxonomy of Generalised Variational Inference.
Notably, our work differs significantly from Bayesian sensitivity analysis, and robustness to the prior of a Bayesian posterior, as discussed for instance in [berger1985, gustafson1995]. We are not concerned with small perturbations of the prior towards some other probability measure, but rather with priors that will in general not result in consistent Bayes posteriors, such as those that are mutually singular with the GVI posterior, or not contain KL neighbourhoods of the data generating process in their support. We provide a fuller account in the appendix.
@@
In this paper, we develop theory concerning GVI posterior measures over infinite dimensional hypothesis spaces $\Theta$. The main contribution of this paper is Theorem 3 where we establish rates of convergence of GVI posterior measures just slower that $n^{-1}$ when using bounded divergences. In doing so we first present sufficient conditions for existence and uniqueness of GVI posteriors, which we then use as a basis to characterise asymptotic concentration to sets containing the minimisers of the loss. Significantly, our theory allows us to deal with instances in which the prior is severely misspecified, such as in the case where the prior has full measure on a subset of $\Theta$ where the Kullback–Leibler divergence rate is infinite and where the prior and GVI posterior are mutually singular. This is highlighted in Figure 1.
@@
This paper is structured as follows. In Section 2 we provide the necessary background for Generalised Variational Inference, prior misspecification, and the assumptions used throughout the paper. Section 3 states and discusses the main results of this paper, Section 4 presents extensions to these, including predictive consistency of GVI under standard losses, and we provide the complete proofs in Section 6. To highlight our results, Section 5 illustrates how GVI is robust to prior misspecification in cases where Bayesian posteriors are inconsistent. In Section 7 we provide some concluding remarks on our results, their relevance to extensions of GVI in Federated Learning [mildner2025], and relevant open problems.
2. Preliminaries
2.1. Notation
@@
Let $X_{1},X_{2},...,X_{n}$ be a sequence of observable random variables, short $X_{1}^{n}$, generated by the abstract probability space $(\Omega,\mathcal{F},P_{0})$ and taking values in the measurable space $(\Xi,\mathcal{X})$ with $P_{0}$ distribution. We consider $n\in\mathbb{N}$ and let this tend to infinity for the concentration results. Denote by $\sigma(X_{1}^{n})$ the natural filtration of this sequence. We do not place any restrictions on the dependence of the data, as long as the loss associated with it satisfies the respective assumptions, especially Assumption 6.
@@
Throughout, we will study probability measures in $\mathcal{P}(\Theta),$ the space of all Borel probability measures on $(\Theta,\mathcal{T})$. Note that $\Theta$ being Polish implies that $\mathcal{P}(\Theta)$ is also Polish, and hence allowing for arbitrary Polish spaces allows for nonparametric inference [kechris1995, dembo2010], we identify $\Theta$ and $\mathcal{P}$}. We begin with some prior $\Pi$ and update this to the GVI posterior $Q_{n}$.
@@
2.2. Assumptions
The assumptions required throughout the paper are primarily related to the behaviour of the data generating process $P_{0}$ and its interaction with the choices of loss functions, divergence, and the set of measures $\mathcal{Q}$}.
@@ 2.2. Assumptions
The loss $L(X_{1}^{n},\,\theta)$ is $\sigma(X_{1}^{n})\times\mathcal{T}$–measurable and lower bounded [^2] Loss functions used in ML are typically non–negative, we, in accordance with the literature on GVI and GBI, allow for more general functions through the lower boundedness. for all $n\in{\mathbb{N}}$.
The next assumption requires for $\mathcal{Q}$} to be a sensible choice which does not make the objective pathological.
@@
In order to reason about the dynamical properties of GVI posteriors we require that minimisers of Equation 1 exist. In particular, we impose the following two assumptions which will be sufficient to prove existence of GVI posteriors in Theorem 1 given that $\mathcal{Q}$} is weak closed.
Assumption 3.
The losses as functions on $\Theta$, $L_{n}:\Theta\rightarrow\mathbb{R}\cup\{+\infty\}$, are coercive and lower semi–continuous for all $n\in\mathbb{N}$.
This assumption is slightly stronger than required as we require for the functional $\nu\mapsto J_{L_{n}}(\nu)$ to be lower semi–continuous; the loss having this property immediately ensures this. As a consequence however, the losses will achieve their infima on any compact set. Moreover, we are able to discard the assumption of the coerciveness of the loss whenever $\Theta$ is compact.
In order for the entire GVI functional $\nu\mapsto T_{n}(\nu)$ to be lower semi–continuous, we require the same to hold for the divergence.
Assumption 4.
The divergence $D:\mathcal{P}(\Theta)\times\mathcal{P}(\Theta)\rightarrow\mathbb{R}_{\geq 0}\cup\{+\infty\}$ is lower semi–continuous in its first argument.
We remark here that this assumption is satisfied by many common choices of divergence including the Kullback–Leibler and the Total Variation. This, together with Assumption 3, provides a key tool in proving the existence and uniqueness of GVI posterior measures in Theorem 1.
To characterise the robustness of GVI to prior misspecification, we analyse the behaviour of GVI posteriors under different priors. In particular, we denote by $\mathcal{G}\subset\mathcal{P}(\Theta)$ the set of possible priors. It turns out that the following assumption provides a sufficient condition for consistency, and in fact rates of convergence, of GVI to hold regardless of the prior $\Pi\in\mathcal{G}$.
@@
It is worth noting that this may be satisfied by both bounded divergences, such as the total variation distance, as well as specific choices of the set of measures we are optimising over, $\mathcal{Q}$, and the set of possible priors, $\mathcal{G}$. In the former case, we may allow for $\mathcal{G}=\mathcal{P}(\Theta)$ and ${\mathcal{Q}}\subset\mathcal{P}(\Theta)$ without restriction. The latter is more restrictive and for instance includes the case where $\mathcal{G}$ is a singleton $\{\Pi\}$ and ${\mathcal{Q}}\subset\{\nu\in\mathcal{P}(\Theta):D(\nu:\Pi)\leq M\}$.
Finally, we require that the dynamics of the data generating process $P_{0}$ and the loss functions, $L_{n}$, mesh sufficiently well together.
@@
This assumption is required for the concentration results of Theorems 2 and 3, and the first part is standard in Bayesian asymptotics (see e.g. [shalizi2009]) where an asymptotic equipartition property, as in the Shannon–MacMillan–Breimann theorem, is required to hold. Unless otherwise stated all convergence takes place in the number of observables $n$.